What is the oxidation state of Cr in the following compound
K2[Cr(CN)2(O)2(O2)(NH3)]
I am confused about (O2). Is it a simple oxygen molecule or peroxide or superoxide?
Please explain.

Grade:

1 Answers

Arun
5 years ago
Dear student
 

K2{Cr(CN)2(O)2(O2)NH3}= k^+2 +{Cr(CN)2(O)2(O2)NH3}^-2

Now oxidation no of

CN(cyano)=-1

O(oxo)=-2

O2(peroxo)=-2

Crcromate)=x

CALCULATING OXIDATION NO OF NEGATIVE PART

x +2(-1)+2(-2)+(-2)=-2

X=2+4+2–2=+6

 

 

Regards

Arun